# Toward Open Vocabulary Aerial Object Detection with CLIP-Activated Student-Teacher Learning

### v2

@@ 4.5 Evaluation on Other Dataset
We further validate our approach on the NWPU VHR-10[6] and DOTA[35]. The specific partitioning methods of the dataset are shown in Table 1. As shown in Table 4, our method achieves 87.6% mAP on NWPU VHR-10 and 58.6% mAP on DOTA.
@@ 5 Conclusion
In this paper, we propose CastDet, a CLIP-activated student-teacher detector designed for open-vocabulary aerial detection. Specifically, we introduce a localization teacher with several reliable box selection strategies (e.g., RPN, BJV, RJV) for novel category discovery. Then, we incorporate RemoteCLIP to acquire limited yet reliable knowledge, serving as an external incentive for student-teacher interactive self-learning. We further introduce a dynamic label queue to store and dynamically update information to obtain richer and more accurate external knowledge. Additionally, we propose a hybrid training approach to simultaneously train multiple data streams, facilitating collaborative training of various sub-modules. With these improvements, CastDet achieves a 40.5% mAP on VisDroneZSD. To the best of our knowledge, this marks the first work on OVD in aerial images. We aspire to lay a foundation for subsequent research in this domain.

### v3

@@ 4.5 Evaluation on Other Dataset
We further validate our approach on the NWPU VHR-10[6] and DOTA[37]. Our method achieves 87.6% mAP on NWPU VHR-10 and 58.6% mAP on DOTA, as shown in Table LABEL:tab:other_dataset_exp.
@@ 5 Conclusion
We propose CastDet, a CLIP-activated student-teacher detector designed for open-vocabulary aerial detection. Specifically, we introduce a localization teacher with several box selection strategies (e.g., RPN, BJV, RJV) for novel category discovery. Then, we incorporate RemoteCLIP to acquire limited yet reliable knowledge, serving as an external incentive for student-teacher interactive self-learning. We further introduce a dynamic label queue to store and dynamically update information to obtain richer and more accurate external knowledge. Additionally, we propose a hybrid training approach to simultaneously train multiple data streams, facilitating collaborative training of various sub-modules. With these improvements, CastDet achieves a 40.5% mAP on VisDroneZSD. To the best of our knowledge, this marks the first work on OVD in aerial images. We aspire to lay a foundation for subsequent research in this domain.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62071333), Fundamental Research Funds for the Central Universities (22120220654), and the Science and Technology on Near-Surface Detection Laboratory (6142414221606).